# TimeBridge: Better Diffusion Prior Design with Bridge Models for Time Series Generation

### v2

@@ Abstract.
Time series generation is widely used in real-world applications such as simulation, data augmentation, and hypothesis testing. Recently, diffusion models have emerged as the de facto approach to time series generation, enabling diverse synthesis scenarios. However, the fixed standard-Gaussian diffusion prior may be ill-suited for general time series data, such as temporal order and fixed points. In this paper, we propose TimeBridge, a framework that flexibly synthesizes time series data by using diffusion bridges to learn paths between a chosen prior and the data distribution. We then explore several prior designs tailored to time series synthesis. Our framework covers (i) data- and time-dependent priors for unconditional generation and (ii) scale-preserving priors for conditional generation. Experiments show that our framework with data-driven priors outperforms standard diffusion models on time series generation.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Modeling methodologiesccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Synthetic time series have expanded their applications by serving roles in data augmentation (Wen et al., 2021), data privacy (Wang et al., 2024), data compression (Zhou et al., 2023), and hypothesis‐driven simulations of data streams (Xia et al., 2024). Common tasks such as imputation and forecasting can be treated as special cases of time series generation. With advances in deep generative modeling, practitioners have developed VAE‐based (Naiman et al., 2024), GAN‐based (Yoon et al., 2019), and diffusion‐based methods (Ho et al., 2020; Karras et al., 2022). Diffusion approaches are especially effective, yielding high‐fidelity and diverse samples (Tashiro et al., 2021; Alcaraz and Strodthoff, 2023) by estimating the reverse trajectory of a forward noising process toward a standard Gaussian prior (Song et al., 2021).
Nevertheless, diffusion models may fail to capture characteristics of time series data because they rely on a fixed mapping from standard Gaussian noise to the data domain. Such restricted priors do not inherently reflect sequential order or scale (Kim et al., 2022), both of which are crucial for modeling historical statistics and temporal dependencies (Zeng et al., 2023). Adding domain‐specific constraints also remains difficult (Harvey et al., 2022; Du et al., 2023); for instance, Coletta et al. (2024) highlighted that generating data under specific bull or bear market patterns requires tailored embedding or penalty functions for sampling guidance.
@@
[…]
Recent studies have revisited the noising process in diffusion models, exploring data-dependent forward and reverse distributions (Han et al., 2022; Yu et al., 2024) and joint consideration of diffusion steps with actual timestamps (Biloš et al., 2023; Chen et al., 2023a). We extend this direction through Schrödinger bridge (Schrödinger, 1932) and diffusion bridges (De Bortoli et al., 2021; Zhou et al., 2024), which learn optimal transport paths between the data distribution and a desired prior. By selecting appropriate fixed priors, diffusion bridges have demonstrated strong performance in conditional applications such as image-to-image translation (Zhou et al., 2024) and text-to-speech (Chen et al., 2023b).
In this paper, we propose TimeBridge, a flexible time series generation framework that addresses multiple time series synthesis with the prior design of diffusion bridge, as shown in Figure 1. Then, we ask what priors are best suited to time series generation as a replacement for the standard Gaussian prior, enabling more efficient and higher-quality synthesis across diverse settings. Against standard diffusion baselines compared in Table 1, the proposed TimeBridge outperforms in both unconditional and conditional tasks, achieving better results across several datasets in Figure 2.
@@ 2. Preliminaries
Consider the multivariate time series ${\mathbf{x}}$ defined within the sample space $\mathcal{X}=\mathbb{R}^{d\times\tau}$, where $d$ represents the dimensionality of the variables and $\tau$ denotes the time length. For each index $i\in\{1,\ldots,N\}$ among the $N$ samples, each time series is denoted as ${\mathbf{x}}^{i}=({\bm{x}}^{i}(1),\ldots,{\bm{x}}^{i}(\tau))\in\mathcal{X}$, given the data ${\bm{x}}^{i}(k)\in\mathbb{R}^{d}$ for each timestamp $k\in\{1,\ldots,\tau\}$. For diffusion steps $t\in\{0,\ldots,T\}$, we denote intermediate time series samples as ${\mathbf{x}}_{t}$, and for a specific instance $i$ as ${\mathbf{x}}_{t}^{i}$.
2.1. Optimal Transport and Schrödinger Bridge
The objective of the Schrödinger Bridge (SB) problem is to find an entropy-regularized optimal transport path between the data distribution $p_{0}$ and a target prior $p_{T}$ by minimizing the KL divergence between paths (Léonard, 2012; Chen et al., 2021):
@@ 2.1. Optimal Transport and Schrödinger Bridge
where, $\mathcal{P}_{[0,T]}$ denotes the space of path measures on the interval $[0,T]$, $p^{\text{ref}}$ is a reference path measure, and $p_{0},p_{T}$ are the marginals of $p$. By choosing a forward stochastic differential equation (SDE) associated with the reference process $p^{\text{ref}}$, the SB can be reformulated as following forward-backward SDEs:
(2) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(3) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are corresponding partial differential equations, detailed in Appendix A.
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940) to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). However, these iterative methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
2.2. Diffusion Bridge
Standard Diffusion Models
In standard diffusion models (Ho et al., 2020), the diffusion process is constructed by gradually injecting noise into samples ${\mathbf{x}}_{0}$ drawn from the data distribution $p_{0}$, forwarding them into a fixed standard Gaussian distribution $p_{T}=\mathcal{N}(\bm{0},\bm{I})$. The corresponding forward SDE is:
(4) $d{\mathbf{x}}_{t}={\mathbf{f}}({\mathbf{x}}_{t},t)dt+g(t)d{\mathbf{w}}_{t},$
where ${\mathbf{f}}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ represents the drift function, $g:[0,T]\rightarrow\mathbb{R}$ denotes the diffusion coefficient, and ${\mathbf{w}}_{t}$ is the Wiener process. As diffusion models learn to reverse the Equation (4), the reverse SDE (Anderson, 1982) is formulated as follows:
where $p({\mathbf{x}}_{t})$ refers to the probability density of ${\mathbf{x}}_{t}$. Diffusion models learn to match the score function $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})$ for denoising (Song et al., 2021).
Diffusion Bridge
The aforementioned standard diffusion models cannot be extended to general prior distributions. Instead, we can construct the diffusion process towards any endpoint by applying Doob’s h-transform (Doob and Doob, 1984) to Equation (4). With a fixed endpoint ${\mathbf{y}}$, the transformed process for the diffusion bridge is as
@@ 2.2. Diffusion Bridge
where ${\mathbf{h}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}$. For a given data distribution $q_{data}({\mathbf{x}},{\mathbf{y}})$, the diffusion bridge (Schrödinger, 1932; De Bortoli et al., 2021) aims to transport ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{T}$ where $({\mathbf{x}}_{0},{\mathbf{x}}_{T})=({\mathbf{x}},{\mathbf{y}})\sim q_{data}({\mathbf{x}},{\mathbf{y}})$ while following above forward process. Similar to the standard diffusion model, we can construct the reverse SDE as follows:
@@
For training, diffusion bridge learns the score function ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$.
Diffusion bridge models have been investigated due to their ability to flexibly map in various domains, including image-to-image translation (Li et al., 2023; Zhou et al., 2024) and text-to-speech synthesis (Chen et al., 2023b), compared to standard diffusion models. Recently, adapting the advances of standard diffusion, such as consistency (He et al., 2024), and deterministic samplers (Zheng et al., 2025), has been further investigated in the diffusion bridge.
@@ 3.1. Scenarios
For time series generation, we cover both unconditional and conditional generations. Unconditional generation considers the total dataset containing $N$ samples, denoted as $\mathcal{D}=\{{{\mathbf{x}}^{i}}\}_{i=1}^{N}$. Our goal is to train a deep generative model to mimic the input dataset and yield synthetic samples statistically similar to a given time series dataset. We aim to make the distribution $p(\hat{\mathbf{x}}|\cdot)$ of synthetic data $\hat{\mathbf{x}}$, similar to the input data distribution $p({\mathbf{x}})$.
Conditional generation further utilizes the paired condition ${\mathbf{y}}$ to guide time series generation of each data sample ${\mathbf{x}}$ as $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. Thus, we try to make the conditional distribution $p(\hat{\mathbf{x}}|{\mathbf{y}})$ of synthetic data $\hat{\mathbf{x}}$ closely mirror the conditional distribution of input data $p({\mathbf{x}}|{\mathbf{y}})$. Considering that the condition in time series is often also a time series as ${\mathbf{y}}\in\mathcal{X}$, Coletta et al. (2024) divided the constraints into two types: (i) soft constraint to guide the time series generation such as trend, or (ii) hard constraint where we should follow during synthesis such as fixed points. In a broader view, the unmasked part in imputation or the historical data in forecasting can be considered as conditions for generation.
@@ 3.2. Diffusion Bridge for Time Series
To approximate the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$ in Equation (8), we adopt the Denoising Diffusion Bridge Model (DDBM) sampler (Zhou et al., 2024). Unlike prior time‐series methods based on the Schrödinger bridge (Chen et al., 2023b; Garg et al., 2024), DDBM aligns with recent diffusion frameworks and enables efficient bridge sampling. Accordingly, we exploit the tractable marginal distribution of ${\mathbf{x}}_{t}$ by setting ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are noise‐schedule functions and ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$. Under the variance‐preserving (VP) schedule, the marginal distribution at time $t$ becomes
@@
For score matching, Karras et al. (2022) proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we match $D_{\theta}$ directly, which is equivalent to training the score function (Karras et al., 2022) as
@@
$:=\frac{{\mathbf{x}}_{t}-(\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})(1-\frac{SNR_{T}}{SNR_{t}}))}{\sigma^{2}_{t}(1-\frac{SNR_{T}}{SNR_{t}})},$
@@
To obtain $D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})$, we follow the backbone of Diffusion-TS (Yuan and Qiao, 2024), which employs an encoder–decoder transformer. Let $w_{(\cdot)}^{i,t}$ denote the input of the interpretable layers for index $i\in{1,\ldots,K}$ at diffusion step $t$. Each decoder block is split into trend and seasonal synthesis layers. First, the trend component is synthesized with a polynomial regressor:
@@
where $X_{tr}^{i,t}$ is the mean value of the output of the $i^{th}$ decoder block with the tensor multiplication ()($\cdot$)( ⋅ ) to $C$.
@@
where argTopK is to get the top $K$ amplitudes and $f_{k}$ represents the Fourier frequency of the index $k$ and $\bar{(\cdot)}$ denotes ()($\cdot$)( ⋅ ) of the conjugates. $A_{i,t}^{(k)}$ and $\Phi_{i,t}^{(k)}$ are the phase and amplitude of the $k$-th frequency, respectively, after the Fourier transform $\mathcal{F}$ as
@@
where $V_{t}^{tr}$ is the output for the trend synthesis layer, $S_{i,t}$ is the output for each seasonal synthesis layer $i$ among $K$ layers and $R$ is the output for estimated residual. Note that calculating Equation (17) is unique in diffusion bridge since the standard diffusion models (Yuan and Qiao, 2024) do not consider the dependency on prior ${\mathbf{x}}_{T}$ for calculating ${\mathbf{x}}_{t}$.
@@
As time series data possess stochasticity (Shen and Kwok, 2023), we directly predict $D_{\theta}$ rather than using the reparametrization technique from noise prediction used in other domains (Karras et al., 2022; Zhou et al., 2024). Moreover, we utilize the second-order Heun sampler (Ascher and Petzold, 1998; Karras et al., 2022), achieving both quality and efficiency over first-order methods.

### v3

@@ Abstract.
Time series generation is widely used in real-world applications such as simulation, data augmentation, and hypothesis testing. Recently, diffusion models have emerged as the de facto approach to time series generation, enabling diverse synthesis scenarios. However, the fixed standard-Gaussian diffusion prior may be ill-suited for time series data, which exhibit properties such as temporal order and fixed time points. In this paper, we propose TimeBridge, a framework that flexibly synthesizes time series data by using diffusion bridges to learn paths between a chosen prior and the data distribution. We then explore several prior designs tailored to time series synthesis. Our framework covers (i) data- and time-dependent priors for unconditional generation and (ii) scale-preserving priors for conditional generation. Experiments show that our framework with data-driven priors outperforms standard diffusion models on time series generation.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780195isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Modeling methodologiesccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series generations have expanded their applications by serving roles in data augmentation (wen2021time), data privacy (wang2024igamt), and hypothesis‐driven simulations of data streams (xia2024market). Even imputation and forecasting can be treated as special cases of time series generation. With advances in deep generative modeling, practitioners have developed VAE‐based (naiman2024generative), GAN‐based (yoon2019time), and diffusion‐based methods (ho2020denoising; karras2022elucidating). Diffusion approaches have shown strong performance, yielding high‐fidelity and diverse samples (tashiro2021csdi; lopez2023diffusionbased) by estimating the reverse trajectory of a forward noising process toward a standard Gaussian prior (song2021scorebased).
Nevertheless, controlling the diffusion process remains difficult (harvey2022flexible; du2023flexible) when incorporating domain-specific constraints or time-dependent properties in time series. For instance, coletta2024constrained highlighted that generating data under specific bull or bear market patterns requires tailored embedding or penalty functions for sampling guidance. As standard diffusion forces a standard Gaussian for prior distribution, it is hard to control the diffusion model to reflect the important properties of time series, such as sequential order, historical statistics, or temporal dependencies (kim2022reversible; zeng2023transformers), during the whole reverse trajectories.
@@
TimeBridge DDBM (zhou2024denoising) 119 ($\downarrow$) $bigcirc \bigcirc \bigcirc \bigcirc$
When viewing a generative model as optimal transport between data and prior distributions, the choice of prior distribution significantly impacts the training difficulty (kollovieh2024flow). Rather than mapping data to a standard normal prior, recent studies altered the diffusion process with data-dependent forward and reverse processes for controlled generation (han2022card; yu2024constructing) or developed diffusion steps aligned with actual timestamps in time series domains (bilovs2023modeling; chen2023provably).
In this paper, we extend this line of work by leveraging diffusion bridges (Schrodinger1932theorie; de2021diffusion; zhou2024denoising) to control prior distributions in time series diffusion models, enabling optimal transport between data and carefully chosen priors. While diffusion bridges have yielded impressive results in conditional generation settings such as image-to-image translation (zhou2024denoising) and text-to-speech (chen2023tts), existing work has focused primarily on how to connect given data and prior distributions, with limited attention to the crucial question of prior selection.
To address this gap in time series generation, we propose TimeBridge, a diffusion bridge model for time series (illustrated in Figure 1) with an architecture tailored to sequential data based on Diffusion-TS (yuan2024diffusionts). We then investigate which priors better replace the standard Gaussian prior, enabling more efficient synthesis and higher-quality results across diverse settings. Compared to standard diffusion baselines summarized in Table 1, TimeBridge shows higher accuracy in both unconditional and conditional tasks across various datasets, as demonstrated in Figure 2.
@@
We publish the code at https://github.com/JinseongP/TimeBridge.
@@ 2. Preliminaries
Consider the multivariate time series $\mathbf{x}$} defined within the sample space $\mathcal{X}=\mathbb{R}^{d\times\tau}$, with the dimensionality $d$ and the time length $\tau$. For each index $i\in\{1,\ldots,N\}$ among the $N$ samples, each time series is denoted as ${\mathbf{x}}^{i}=({\bm{x}}^{i}(1),\ldots,{\bm{x}}^{i}(\tau))\in\mathcal{X}$, given the data ${\bm{x}}^{i}(k)\in\mathbb{R}^{d}$ for each timestamp $k\in\{1,\ldots,\tau\}$. For diffusion steps $t\in\{0,\ldots,T\}$, we denote intermediate time series samples as ${\mathbf{x}}_{t}$, and for a specific instance $i$ as ${\mathbf{x}}_{t}^{i}$.
Diffusion Models.
In standard diffusion models (ho2020denoising), the diffusion process is constructed by gradually injecting noise into samples ${\mathbf{x}}_{0}$ drawn from the data distribution $p_{0}$, forwarding them into a fixed standard Gaussian distribution $p_{T}=\mathcal{N}(\bm{0},\bm{I})$. The corresponding forward SDE is:
(1) $d{\mathbf{x}}_{t}={\mathbf{f}}({\mathbf{x}}_{t},t)dt+g(t)d{\mathbf{w}}_{t},$
where ${\mathbf{f}}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ represents the drift function, $g:[0,T]\rightarrow\mathbb{R}$ denotes the diffusion coefficient, and ${\mathbf{w}}_{t}$ is the Wiener process. As diffusion models learn to reverse the Equation (1), the reverse SDE (anderson1982reverse) is formulated as follows:
where $p({\mathbf{x}}_{t})$ refers to the probability density of ${\mathbf{x}}_{t}$. Diffusion models learn to match the score function $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})$ for denoising (song2021scorebased).
Optimal Transport and Schrödinger Bridge.
The aforementioned standard diffusion models assume a standard Gaussian in the prior, and thus cannot be extended to general prior distributions. On the other hand, the optimal transport view for generative modeling is to find a trajectory of two arbitrary distributions. Specifically, the Schrödinger Bridge (SB) provides an entropy-regularized solution by minimizing the KL divergence between path measures (chen2021likelihood):
@@
where, $\mathcal{P}_{[0,T]}$ denotes the space of path measures on the interval $[0,T]$, $p^{\text{ref}}$ is a reference path measure, and $p_{0},p_{T}$ are the marginals of $p$. By choosing a forward stochastic differential equation (SDE) associated with the reference process $p^{\text{ref}}$, the SB can be reformulated by the following forward-backward SDEs:
(4) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\ {\mathbf{x}}_{0}\sim p_{data},\\ d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\ {\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are corresponding partial differential equations. In earlier practice, the Schrödinger bridge-based trajectory simulation typically relied on iterative methods (kullback1968probability), making it less widely used than diffusion models due to computational complexity and training stability (zhou2024denoising). The details of Schrödinger Bridge, including iterative solution, are explained in Appendix A.
Diffusion Bridge.
Recently, a non-iterative approach to diffusion bridges has emerged, which constructs diffusion processes towards any endpoint by applying Doob’s h-transform (doob1984classical) to Equation (1). With a fixed endpoint $\mathbf{y}$}, the objective is to trace the trajectory of the transformed path for the diffusion bridge as follows:
@@
where ${\mathbf{h}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}$. For a given data distribution $q_{data}({\mathbf{x}},{\mathbf{y}})$, the diffusion bridge (Schrodinger1932theorie; de2021diffusion) aims to transport ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{T}$ where $({\mathbf{x}}_{0},{\mathbf{x}}_{T})=({\mathbf{x}},{\mathbf{y}})\sim q_{data}({\mathbf{x}},{\mathbf{y}})$ while following above forward process. Similar to the standard diffusion model, we can construct the reverse SDE as follows:
@@
For training, diffusion bridge can leverage score-based methods by learning the score function ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$.
Diffusion bridge models have been investigated due to their ability to flexibly map in various domains, including image-to-image translation (li2023bbdm; zhou2024denoising) and text-to-speech synthesis (chen2023tts), compared to standard diffusion models. Recently, adapting the advances of standard diffusion, such as consistency (he2024consistency), and deterministic samplers (zheng2025diffusion), has been further investigated in the diffusion bridge.
@@ 3.1. Scenarios
[…]
Conditional generation further utilizes the paired condition $\mathbf{y}$} to guide time series generation of each data sample $\mathbf{x}$} as $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. Thus, we try to make the conditional distribution $p(\hat{\mathbf{x}}|{\mathbf{y}})$ of synthetic data $\hat{\mathbf{x}}$ closely mirror the conditional distribution of input data $p({\mathbf{x}}|{\mathbf{y}})$. Considering that the condition in time series is often also a time series as ${\mathbf{y}}\in\mathcal{X}$, coletta2024constrained divided the constraints into two types: (i) soft constraint to guide the time series generation such as trend, or (ii) hard constraint where we should follow during synthesis such as fixed points. In a broader view, the unmasked part in imputation or the historical data in forecasting can be considered as conditions for generation.
(a) Unconditional
(b) Linear trend
(c) Polynomial trend
(d) Fixed point
(e) Missing value
@@ 3.2. Diffusion Bridge for Time Series
To approximate the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$ in Equation (7), we adopt the Denoising Diffusion Bridge Model (DDBM) sampler (zhou2024denoising). Unlike prior time‐series methods based on the Schrödinger bridge (chen2023tts; garg2024soft), DDBM aligns with recent diffusion frameworks and enables efficient bridge sampling. Accordingly, we exploit the tractable marginal distribution of ${\mathbf{x}}_{t}$ by setting ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are noise‐schedule functions and ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$. Under the variance‐preserving (VP) schedule, the marginal distribution at time $t$ becomes
@@
For score matching, karras2022elucidating proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we match $D_{\theta}$ directly, which is equivalent to training the score function (karras2022elucidating) as
@@
$=\frac{{\mathbf{x}}_{t}-(\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})(1-\frac{SNR_{T}}{SNR_{t}}))}{\sigma^{2}_{t}(1-\frac{SNR_{T}}{SNR_{t}})},$
@@
To obtain $D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})$, we follow the backbone of Diffusion-TS (yuan2024diffusionts), which employs an encoder–decoder transformer. Let $w_{(\cdot)}^{i,t}$ denote the input of the interpretable layers for index $i\in{1,\ldots,K}$ at diffusion step $t$. Each decoder block is split into trend and seasonal synthesis layers. First, the trend component is synthesized with a polynomial regressor:
@@
where $X_{tr}^{i,t}$ is the mean value of the output of the $i^{th}$ decoder block with the tensor multiplication $(\cdot)$ to $C$.
@@
where argTopK is to get the top $K$ amplitudes and $f_{k}$ represents the Fourier frequency of the index $k$ and $\bar{(\cdot)}$ denotes $(\cdot)$ of the conjugates. $A_{i,t}^{(k)}$ and $\Phi_{i,t}^{(k)}$ are the phase and amplitude of the $k$-th frequency, respectively, after the Fourier transform $\mathcal{F}$ as
@@
where $V_{t}^{tr}$ is the output for the trend synthesis layer, $S_{i,t}$ is the output for each seasonal synthesis layer $i$ among $K$ layers and $R$ is the output for estimated residual. Note that calculating Equation (16) is unique in diffusion bridge since the standard diffusion models (yuan2024diffusionts) do not consider the dependency on prior ${\mathbf{x}}_{T}$ for calculating ${\mathbf{x}}_{t}$.
@@
As time series data possess stochasticity (shen2023non), we directly predict $D_{\theta}$ rather than using the reparametrization technique from noise prediction used in other domains (karras2022elucidating; zhou2024denoising). Moreover, we utilize the second-order Heun sampler (ascher1998computer; karras2022elucidating), achieving both quality and efficiency over first-order methods.